OpenStudy (anonymous):

a rectangle's width is twice as long as its length if the perimeter is 36 in. find its length A 4in B 6 in C 8 in D 10 in

OpenStudy (anonymous):

That question doesn't make sense. Are you sure you wrote it right?

OpenStudy (anonymous):

and its a rectangles width and meant to put use an algebraic expression

OpenStudy (anonymous):

w= width l=length So we know that 2l=w and we know that 2w+2l=36

OpenStudy (anonymous):

So we can substitute 2l in for w in the second equation like so 2(2l)+2l=36 4l+2l=36 6l=36 l=6 and w=12 So 2(12) + 2(6)=36 24+12+36 36=36 So length is 6

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

No problem.